# BalDRO: A Distributionally Robust Optimization based Framework for Large Language Model Unlearning

### v1

@@ Abstract.
As Large Language Models (LLMs) increasingly shape online content, how to remove targeted information from well-trained LLMs (also known as LLM unlearning) has become increasingly critical for web governance. A key challenge in LLM unlearning lies in the sample-wise imbalance within the forget set: different samples exhibit widely varying unlearning difficulty, leading to asynchronous forgetting speeds where some knowledge remains insufficiently erased while others become over-forgotten. To address this challenge, we propose BalDRO, a novel and efficient framework for balanced LLM unlearning. BalDRO formulates unlearning as a min–sup process, where the inner process identifies a worst-case data distribution that adaptively emphasizes hard-to-unlearn samples, while the outer process updates model parameters based on the worst-case data distribution. We instantiate this formulation through two efficient variants: BalDRO-G, a discrete GroupDRO-based approximation that focuses on high-loss subsets, and BalDRO-DV, a continuous Donsker–Varadhan dual method that enables smooth, adaptive weighting within standard LLM training pipelines. Extensive experiments on the TOFU and MUSE benchmarks demonstrate the effectiveness of our proposed BalDRO, yielding significant improvements in both forgetting quality and model utility over existing methods. For reproducibility, we have anonymously released the code for BalDRO [^1] https://github.com/nxZhai/BalDRO.
@@
This is a preprint. The final version will appear in The Web Conference 2026.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
As Large Language Models (LLMs) become increasingly embedded in web platforms and services (deng2024large; levy2024st; yan2025agentsociety), ensuring that these models behave in a trustworthy and responsible manner has become essential for maintaining the reliability of web-based information ecosystems. A key aspect of achieving such reliability is the ability to remove outdated, incorrect, or privacy-sensitive knowledge from LLMs so that their behavior remains aligned with user expectations and public values (geng2025comprehensive; ye2025towards). This need is further reinforced by legal frameworks such as the General Data Protection Regulation (GDPR) [^2] https://gdpr-info.eu/ and the California Consumer Privacy Act (CCPA) [^3] https://oag.ca.gov/privacy/ccpa , which mandate the “right to be forgotten” and require machine learning systems to support verifiable data erasure. Together, these factors highlight the importance of developing effective LLM unlearning techniques (zhao2024makes; chen2025safeeraser).
@@
Among the major challenges in achieving effective LLM unlearning, a particularly critical one lies in the heterogeneous unlearning difficulty across samples in the forget set (krishnan2025not; fan2024simplicity; zhangnegative). As illustrated in Figure 1(a), when applying NPO (zhangnegative) to the same forget set, samples that start from similar initial states quickly diverge in their perplexity (PPL) trajectories at the same training epoch. This divergence reveals that different samples are forgotten at substantially different rates—some being easy to erase, while others remain resistant. Figure 1(b) further highlights this imbalance: each sample reaches its convergence point at a different epoch, making it difficult to determine when the entire forget set has been properly unlearned. This asynchronous forgetting dynamic is particularly problematic for gradient-based LLM unlearning methods, as such variants of the negative cross-entropy loss lack a well-defined upper bound (liu2025rethinking). Continuing optimization to ensure full erasure of hard samples inevitably over-forgets the easy ones, thereby degrading overall model utility. This phenomenon motivates a central question: how can we achieve a balanced state where all forget samples are unlearned simultaneously?
@@
  • We propose BalDRO, a novel and efficient DRO-based framework that balances sample during unlearning. Specifically, BalDRO  formulates LLM unlearning as a min–sup bi-level process, and provides two tractable realizations for the inner process.
  • We conduct extensive experiments on TOFU and MUSE benchmarks, demonstrating the effectiveness of BalDRO  on achieving a better tradeoff between forgetting quality and model utility. For example, on the TOFU benchmark, BalDRO-G and BalDRO-DV based on NPO improve forget quality by more than 20% over NPO, while also delivering modest gains in model utility.
@@ 2.1. LLM Unlearning
LLM unlearning aims to suppress knowledge contained in the forget set while preserving performance on the retain set (liu2025rethinking; jia2024soul; jeung2025dusk; liu2024large), encompassing both model-editing–based approaches (shen2025llm; yu2025unierase; li2025cross; li2025editing) and gradient-based approaches (mekala2025alternate; fan2024simplicity; yangexploring). Among them, we focus on gradient-based methods, as they are model-agnostic and compatible with LLM finetuning pipelines. Gradient-based methods can be broadly grouped into two categories. The first line of work is targeted unlearning, which defines explicit substitute outputs—typically refusal-style responses—for each forget-set query. These methods treat refusals as positive examples and enforce them using preference-based objectives such as DPO (rafailov2023direct). Subsequent variants expand this paradigm: FLAT incorporates $f$-divergence–based loss adjustments (wangllm), while AltPO generates diverse positive alternatives by substituting knowledge-relevant tokens (mekala2025alternate). However, targeted methods may induce shortcut behaviors (qisafety), where the model learns to mimic refusal templates based on patterns in the input.
@@ 3. Preliminary
(2) $\ell^{\mathrm{GA}}_{f}(\theta)=-\mathbb{E}_{(x,y)\in{D}_{f}}\!\left[\ell_{f}\!\left(y\mid x;\theta\right)\right].$
@@ 4.2.1. BalDRO-G: Discrete Group-Level Approximation
Intuitively, this formulation serves as a discrete instantiation of the inner process in Definition 4.1: the adversarial distribution $Q_{f}$ is approximated by assigning all probability mass to the group with the largest expected forget loss. This allows BalDRO-G to operationalize the theoretical min–sup objective through a finite, group-level approximation that is both stable and computationally efficient in practice. At each iteration, BalDRO-G enforces progress on the worst-performing group, directing the optimization toward the under-forgotten region of the data space. Rather than relying on predefined grouping schemes, we adopt a percentile-based strategy (top-50%) to avoid the instability of extreme single-sample maxima while still focusing on the hardest region of the distribution. This design follows standard practice in robust optimization and is further validated in Appendix B.
@@ 4.3.1. Additional Time Complexity
In summary, both BalDRO-G and BalDRO-DV preserve the scalability of existing unlearning pipelines, adding only minimal constant-time operations while maintaining the same asymptotic complexity.
@@ 5. Experiments
Due to page limits, we have placed additional experiments in Appendix B, including varying the sample selection ratio in BalDRO-G.
@@ 5.1.2. Evaluation Metrics
We compare our method with several baselines. Original refers to the model trained on both forget and retain sets; it represents the starting point of unlearning. In contrast, Retain denotes the model fine-tuned only on the retain set, reflecting an idealized upper bound for LLM unlearning. The remaining baselines, introduced in Section 3, cover the classic and recent gradient-based unlearning methods. Specifically, we include GradAscent (GA) (yao2024large), which performs gradient ascent on forget samples to reduce the model’s confidence, and GradDiff (GD) (yao2024large), which additionally leverages the retain set to mitigate unnecessary utility loss. Moreover, we evaluate BalDRO on three representative and state-of-the-art unlearning frameworks—NPO (zhangnegative), SimNPO (fan2024simplicity), and SatImp (yangexploring). These baselines constitute the latest gradient-based unlearning techniques, while editing-based approaches are not included (shen2025llm; yu2025unierase), as they follow a fundamentally different model-editing paradigm that is incompatible with us.
@@
All experiments are conducted on a server equipped with 8 NVIDIA A800 GPUs. We use LLaMA-2-7B (touvron2023llama) as the primary backbone model for evaluating overall unlearning performance. For all methods, we perform a hyperparameter search over learning rates in $\{1\times 10^{-5},2\times 10^{-5},5\times 10^{-5},1\times 10^{-4}\}$, batch sizes in $\{8,16,32\}$, $\beta$ in $\{1.0,2.0,5.0,10.0\}$ and $\lambda$ in $\{0.25,0.5,1.0,2.0\}$. This setup ensures a fair and consistent comparison across methods, allowing the reported performance to reflect the best achievable results under each experimental configuration. More detailed information can be found in our released code [^4] https://anonymous.4open.science/r/anonymous_web4good2026_id480-2DE5.
Figure 3. Performance with varying forget ratios (5% and 10%) on the TOFU benchmark. We focus on FQ and MU, the two most commonly used metrics on TOFU, and select NPO and SimNPO as base methods due to their strong overall performance on these two metrics. Here, “+G” denotes our proposed BalDRO-G, and “+DV” corresponds to BalDRO-DV. The dashed horizontal line indicates the retain baseline.
@@ 5.2.1. Main Results on TOFU
Table 1 presents the results under the TOFU benchmark. We summarize three observations. First, BalDRO variants consistently improve FQ and reduce EM/ES across most base methods, showing that balancing the forgetting rates of samples in the forget set effectively strengthens the unlearning process. Second, different base methods benefit to different degrees. SimNPO obtains the most stable overall gains: FQ rises from 0.4046 to 0.5786, EM decreases, and several TR metrics also improve. In contrast, SatImp shows the largest relative jump in FQ (0.0013→0.0971 (BalDRO-G), over 70×), mainly because its baseline forgetting is extremely weak, leaving enough space for improvements. Third, BalDRO-DV and BalDRO-G show different strengths across backbones. For NPO and SimNPO, BalDRO-DV typically performs better, achieving the highest FQ and lower EM/ES while maintaining competitive TR performance. For SatImp, however, BalDRO-G works better, giving both the strongest FQ improvement and more stable TR results.
@@ 5.2.2. Main Results on MUSE
Table 2 summarizes the results on the MUSE benchmark. We have several observations from this table. First, we observe that both BalDRO-G and BalDRO-DV consistently enhance the base methods (NPO, SimNPO, SatImp) across domains. This demonstrates the general effectiveness of our framework: BalDRO  improves forget quality while maintaining LLM utility. For instance, on the News domain, BalDRO-G based on NPO lowers VM-Df (0.4255→0.3826) and improves PL (–90.85→–65.70), indicating better forget quality and better model utility. Second, across both domains, BalDRO-G and BalDRO-DV perform comparably. Rather than one variant uniformly dominating the other, they exhibit complementary strengths across settings. Given that BalDRO-DV only improves the objective function and is therefore simpler to implement, it may be the more practical choice in real-world deployments. Finally, the three base methods (NPO, SimNPO, SatImp) respond to BalDRO  with different degrees of improvement. NPO already delivers strong baseline, limiting the potential gains. E.g., in the Books domain, BalDRO-G yields modest changes in KM-Dr (0.6424→0.6486) and PL (–55.77→–54.22). In contrast, SimNPO and SatImp perform worse, giving BalDRO  substantially more room to provide larger improvements across different metrics.
@@ 5.3.4. Performance on Other Metrics (RQ 5)
We have several findings from Table 3. First, “Original” obtains values close to 1.0 across all four metrics, indicating substantial memorization and thus the highest privacy risk. Second, all unlearning baselines substantially reduce these scores, confirming that they effectively mitigate membership inference attacks. Third, introducing BalDRO  yields consistent improvements across base methods. For instance, within NPO, BalDRO-DV lowers LOSS from 0.4681 to 0.3481 and ZLib from 0.4594 to 0.3250, showing clearer privacy gains. BalDRO-G also improves privacy, albeit to a slightly lesser degree, reflecting BalDRO-DV’s stronger robustness in this setting. The most pronounced improvements occur when BalDRO  is combined with SimNPO. LOSS drops from 0.4925 to 0.1769 and MinK++ from 0.2738 to 0.1075—representing the best overall performance across all configurations. This demonstrates a strong synergy: BalDRO  effectively amplifies SimNPO’s forgetting signal and further suppresses residual memorization. In contrast, SatImp benefits to a more limited extent, suggesting that the impact of BalDRO  may depend on the characteristics of the base method. Nevertheless, across all cases, BalDRO  consistently contributes to measurable reductions in privacy risk, reinforcing its effectiveness as a general enhancement to LLM unlearning.
@@
Table 3. Extended metrics evaluated on different methods on the TOFU benchmark with forget ratio = 1%. We bold the best result.
@@ 6. Conclusion and Future Work
How to unlearn specific information from large language models is essential for responsible and trustworthy web-scale AI. A key challenge lies in the sample-wise imbalance within the forget set: easy samples disappear quickly under unlearning updates, but harder samples retain their influence for much longer, causing the overall forgetting process to become highly uneven. To address this issue, we viewed LLM unlearning through distributional robustness and developed BalDRO, a unified framework that adaptively balances sample contributions during unlearning. We then proposed its two realizations (BalDRO-G and BalDRO-DV), which provide discrete and continuous mechanisms for achieving synchronized forgetting across samples. We conducted extensive experiments on the TOFU and MUSE benchmarks, to demonstrate that BalDRO  consistently improves both forget quality and model utility compared to existing unlearning methods. In the future, we plan to further investigate the balance between forget quality and general utility by redesigning the loss function to reduce cost of LLM unlearning.

### v2

@@ Abstract.
As Large Language Models (LLMs) increasingly shape online content, how to remove targeted information from well-trained LLMs (also known as LLM unlearning) has become increasingly critical for web governance. A key challenge in LLM unlearning lies in the sample-wise imbalance within the forget set: different samples exhibit widely varying unlearning difficulty, leading to asynchronous forgetting speeds where some knowledge remains insufficiently erased while others become over-forgotten. To address this challenge, we propose BalDRO, a novel and efficient framework for balanced LLM unlearning. BalDRO formulates unlearning as a min–sup process, where the inner process identifies a worst-case data distribution that adaptively emphasizes hard-to-unlearn samples, while the outer process updates model parameters based on the worst-case data distribution. We instantiate this formulation through two efficient variants: BalDRO-G, a discrete GroupDRO-based approximation that focuses on high-loss subsets, and BalDRO-DV, a continuous Donsker–Varadhan dual method that enables smooth, adaptive weighting within standard LLM training pipelines. Extensive experiments on the TOFU and MUSE benchmarks demonstrate the effectiveness of our proposed BalDRO, yielding significant improvements in both forgetting quality and model utility over existing methods. For reproducibility, we have released the code for BalDRO [^1] https://github.com/nxZhai/BalDRO.
@@
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792975isbn: 979-8-4007-2307-0/2026/04ccs: Security and privacy Privacy protectionsccs: Computing methodologies Natural language processingccs: Computing methodologies Machine learning
@@ 1. Introduction
As Large Language Models (LLMs) become increasingly embedded in web platforms and services (liu2025debate; yan2025agentsociety; bai2025learning; hu2024psycollm; xu2025multiagentesc), ensuring that these models behave in a trustworthy and responsible manner has become essential for maintaining the reliability of web-based information ecosystems (qin2025explainable; xia2025identifying; zhou2025semi; zhou2025logic). A key aspect of achieving such reliability is the ability to remove outdated, incorrect, or privacy-sensitive knowledge from LLMs so that their behavior remains aligned with public values (geng2025comprehensive; ye2025towards; xu2025multi), and safety requirements (ma2025attackseqbench; fang2025we). This need is further reinforced by legal frameworks such as the General Data Protection Regulation (GDPR) and the California Consumer Privacy Act (CCPA) , which mandate the “right to be forgotten” and require machine learning systems to support verifiable data erasure (zhao2024advanchor; zhao2024separable). Together, these factors highlight the importance of developing effective LLM unlearning techniques (zhao2024makes; chen2025safeeraser).
@@
Among the major challenges in achieving effective LLM unlearning, a particularly critical one lies in the heterogeneous unlearning difficulty across samples in the forget set (krishnan2025not; fan2024simplicity; zhangnegative). Similar sample-level difficulty heterogeneity has also been observed in conditional generative models (e.g., progressive diffusion and unified conditional person generation frameworks) also exhibit strong heterogeneity across conditioning patterns, leading to uneven optimization difficulty (shen2024advancing; shen2024imagpose). As illustrated in Figure 1(a), when applying NPO (zhangnegative) to the same forget set, samples that start from similar initial states quickly diverge in their perplexity (PPL) trajectories at the same training epoch. This divergence reveals that different samples are forgotten at substantially different rates—some being easy to erase, while others remain resistant. Figure 1(b) further highlights this imbalance: each sample reaches its convergence point at a different epoch, making it difficult to determine when the entire forget set has been properly unlearned. This asynchronous forgetting dynamic is particularly problematic for gradient-based LLM unlearning methods, as such variants of the negative cross-entropy loss lack a well-defined upper bound (liu2025rethinking). Continuing optimization to ensure full erasure of hard samples inevitably over-forgets the easy ones, thereby degrading overall model utility. This phenomenon motivates a central question: how can we achieve a balanced state where all forget samples are unlearned simultaneously?
@@
  • We propose BalDRO, a novel and efficient DRO-based framework that balances samples. Specifically, BalDRO  formulates LLM unlearning as a min–sup bi-level process, and provides two tractable realizations for the inner process.
  • We conduct extensive experiments on TOFU and MUSE benchmarks, demonstrating the effectiveness of BalDRO  on achieving a better tradeoff between forgetting quality and model utility, e.g., on the TOFU benchmark, BalDRO-G and BalDRO-DV based on NPO improve forget quality by more than 20% over NPO, while also delivering modest gains in model utility.
@@ 2.1. LLM Unlearning
LLM unlearning aims to suppress knowledge contained in the forget set while preserving performance on the retain set (liu2025rethinking; jia2024soul; jeung2025dusk; liu2024large; zhai2026maximizing), encompassing both model-editing–based approaches (shen2025llm; yu2025unierase; pan2025precise; li2025editing; pan2024finding) and gradient-based approaches (mekala2025alternate; fan2024simplicity; yangexploring). Among them, we focus on gradient-based methods, as they are model-agnostic and compatible with LLM finetuning pipelines. Gradient-based methods can be broadly grouped into two categories. The first line of work is targeted unlearning, which defines explicit substitute outputs—typically refusal-style responses—for each forget-set query. These methods treat refusals as positive examples and enforce them using preference-based objectives such as DPO (rafailov2023direct). Subsequent variants expand this paradigm: FLAT incorporates $f$-divergence–based loss adjustments (wangllm), while AltPO generates diverse positive alternatives by substituting knowledge-relevant tokens (mekala2025alternate). However, targeted methods may induce shortcut behaviors (qisafety), where the model learns to mimic refusal templates based on patterns in the input.
@@ 3. Preliminary
(2) $\mathcal{L}^{\mathrm{GA}}_{f}(\theta)=\mathbb{E}_{(x,y)\sim D_{f}}\big[\log\pi_{\theta}(y\mid x)\big].$
@@ 4.2.1. BalDRO-G: Discrete Group-Level Approximation
Intuitively, this formulation serves as a discrete instantiation of the inner process in Definition 4.1: the adversarial distribution $Q_{f}$ is approximated by assigning all probability mass to the group with the largest expected forget loss. This allows BalDRO-G to operationalize the theoretical min–sup objective through a finite, group-level approximation that is both stable and computationally efficient in practice. At each iteration, BalDRO-G enforces progress on the worst-performing group, directing the optimization toward the under-forgotten region of the data space. Rather than relying on predefined grouping schemes, we adopt a percentile-based strategy (top-50%) to avoid the instability of extreme single-sample maxima while still focusing on the hardest region of the distribution.
@@ 5.1.2. Evaluation Metrics
We compare our method with several baselines. Original refers to the model trained on both forget and retain sets; it represents the starting point of unlearning. In contrast, Retain denotes the model fine-tuned only on the retain set, reflecting an idealized upper bound for LLM unlearning. The remaining baselines, introduced in Section 3, cover the classic and recent gradient-based unlearning methods. Specifically, we include GradAscent (GA) (yao2024large), which performs gradient ascent on forget samples to reduce the model’s confidence, and GradDiff (GD) (yao2024large), which additionally leverages the retain set to mitigate unnecessary utility loss. Moreover, we evaluate BalDRO on three representative and state-of-the-art unlearning frameworks—NPO (zhangnegative), SimNPO (fan2024simplicity), and SatImp (yangexploring). These baselines constitute the latest gradient-based unlearning techniques, while editing-based approaches are not included (shen2025llm; yu2025unierase; zeng2025visual), as they follow a fundamentally different model-editing paradigm that is incompatible with us.
@@
All experiments are conducted on a server equipped with 8 NVIDIA A800 GPUs. We use LLaMA-2-7B (touvron2023llama) as the primary backbone model for evaluating overall unlearning performance. For all methods, we perform a hyperparameter search over learning rates in $\{1\times 10^{-5},2\times 10^{-5},5\times 10^{-5},1\times 10^{-4}\}$, batch sizes in $\{8,16,32\}$, $\beta$ in $\{1.0,2.0,5.0,10.0\}$ and $\lambda$ in $\{0.25,0.5,1.0,2.0\}$.
Figure 3. Performance with varying forget ratios (5% and 10%) on the TOFU benchmark. We focus on FQ and MU, the two most commonly used metrics on TOFU, and select NPO and SimNPO as base methods due to their strong overall performance on these two metrics. Here, “+G” denotes our proposed BalDRO-G, and “+DV” corresponds to BalDRO-DV.
@@ 5.2.1. Main Results on TOFU
Table 1 presents the results under the TOFU benchmark. We summarize three observations. First, BalDRO variants consistently improve FQ and reduce EM/ES across most base methods. Second, different base methods benefit to different degrees. SimNPO obtains the most stable overall gains: FQ rises from 0.4046 to 0.5786, EM decreases, and several TR metrics also improve. In contrast, SatImp shows the largest relative jump in FQ (0.0013→0.0971 (BalDRO-G), over 70×), mainly because its baseline forgetting is extremely weak, leaving enough space for improvements. Third, BalDRO-DV and BalDRO-G show different strengths across backbones. For NPO and SimNPO, BalDRO-DV typically performs better, achieving the highest FQ and lower EM/ES while maintaining competitive TR performance. For SatImp, however, BalDRO-G works better, giving both the strongest FQ improvement and more stable TR results.
@@ 5.2.2. Main Results on MUSE
Table 2 summarizes the results on the MUSE benchmark. We have several observations from this table. First, we observe that both BalDRO-G and BalDRO-DV consistently enhance the base methods (NPO, SimNPO, SatImp) across domains. This demonstrates the general effectiveness of our framework: BalDRO  improves forget quality while maintaining LLM utility. For instance, on the News domain, BalDRO-G based on NPO lowers VM-Df (0.4255→0.3826) and improves PL (–90.85→–65.70). Second, across both domains, BalDRO-G and BalDRO-DV perform comparably. Rather than one variant uniformly dominating the other, they exhibit complementary strengths across settings. Given that BalDRO-DV only improves the objective function and is therefore simpler to implement, it may be the more practical choice in real-world deployments. Finally, the three base methods (NPO, SimNPO, SatImp) respond to BalDRO  with different degrees of improvement. NPO already delivers strong baseline, limiting the potential gains. E.g., in the Books domain, BalDRO-G yields modest changes in KM-Dr (0.6424→0.6486) and PL (–55.77→–54.22). In contrast, SimNPO and SatImp perform worse, giving BalDRO  substantially more room to provide larger improvements across different metrics.
@@ 5.3.4. Performance on Other Metrics (RQ 5)
We have several findings from Table 3. First, “Original” obtains values close to 1.0 across all four metrics, indicating substantial memorization and thus the highest privacy risk. Second, all unlearning baselines substantially reduce these scores, confirming that they effectively mitigate membership inference attacks. Third, introducing BalDRO  yields consistent improvements across base methods. For instance, within NPO, BalDRO-DV lowers LOSS from 0.4681 to 0.3481 and ZLib from 0.4594 to 0.3250, showing clearer privacy gains. BalDRO-G also improves privacy, albeit to a slightly lesser degree, reflecting BalDRO-DV’s stronger robustness in this setting. The most pronounced improvements occur when BalDRO  is combined with SimNPO. This demonstrates a strong synergy: BalDRO  effectively amplifies SimNPO’s forgetting signal and further suppresses residual memorization.
@@
Table 3. Extended metrics evaluated on different methods on the TOFU benchmark with forget ratio = 1%.
@@ 6. Conclusion and Future Work
How to unlearn specific information from LLMs is essential for trustworthy web-scale AI. A key challenge lies in the sample-wise imbalance within the forget set: easy samples disappear quickly under unlearning updates, but harder samples retain their influence for much longer. To address this issue, we viewed LLM unlearning through distributional robustness and developed BalDRO, a unified framework that adaptively balances sample contributions during unlearning. We then proposed its two realizations (BalDRO-G and BalDRO-DV), which provide discrete and continuous mechanisms for achieving synchronized forgetting across samples. We conducted extensive experiments on the TOFU and MUSE benchmarks, to demonstrate that BalDRO  consistently improves both forget quality and model utility compared to existing unlearning methods. In the future, we plan to further investigate the balance between forget quality and general utility by redesigning the loss function to reduce cost of LLM unlearning.
Acknowledgements.
This work has been supported by grants from the National Natural Science Foundation of China under Grant 72188101 and Grant U22A2094.